OpenStudy (anonymous):

Solve for y: 3x^2 - 7xy - 6y2 = 4 using quadratic formula...

OpenStudy (anonymous):

help me po with this problem .. i don't know what to do .. huhu T_T

OpenStudy (anonymous):

@terenzreignz please help me po. :(

terenzreignz (terenzreignz):

Do you know the quadratic formula? \[\Large \frac{-b\pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

yes po . :)

terenzreignz (terenzreignz):

No using po. I'm not gurang :P

terenzreignz (terenzreignz):

Anyway, I shall rewrite your equation... \[\Large -6y^2 -7xy +3x^2=0\]

OpenStudy (anonymous):

HAHAHA . .okay .. kung yan ang gusto. HEHE.

terenzreignz (terenzreignz):

Now, can you solve for y?

OpenStudy (anonymous):

wait.. where is the 4??

terenzreignz (terenzreignz):

Whoops...

terenzreignz (terenzreignz):

\[\Large -6y^2 -7xy +3x^2=4\]

terenzreignz (terenzreignz):

We need 0 on the right-hand-side, so minus 4, on both sides... \[\Large -6y^2 -7xy +3x^2-4=0\]

OpenStudy (anonymous):

ahh.. yes.. then, how will I know the (a, b, and c) ?

terenzreignz (terenzreignz):

Now, here's subtle hint... \[\Large \color{red}ay^2 +\color{blue}by+\color{green}c=0\]

OpenStudy (anonymous):

ahh.. then ??

terenzreignz (terenzreignz):

\[\Large \color{red}{(-6)}y^2\color{blue}{-7x}y+\color{green}{(3x^2 - 4)}=0\] Does that give you a clue? :P

OpenStudy (anonymous):

ahh.. yes .. :) then , i will substitute that in quadratic formula na?

terenzreignz (terenzreignz):

Yup. Easy as pie...

OpenStudy (anonymous):

ahh .. e,  .. ganyan ?